OpenStudy (anonymous):

Which point lies on the graph of the following piecewise function? y = x - 1 if x < 2 y = x2 if x greater than or equal to 2

OpenStudy (anonymous):

(-1, 1) (1, 1) (2, 1) (3, 9)

OpenStudy (anonymous):

This is the graph

OpenStudy (anonymous):

I know that the open circles mean that it is less than the number and the open circles are equal to or to the number

ganeshie8 (ganeshie8):



OpenStudy (anonymous):

So (3,9) by counting the rise over run?

ganeshie8 (ganeshie8):

Look at the given function ``` y = x - 1 if x < 2 y = x^2 if x greater than or equal to 2 ```

ganeshie8 (ganeshie8):

try each option, the first option is (-1, 1) since the x value, -1 is less than 2, you need to use first case for working y : y = -1-1 = -2 so (-1, -2) is a point on the graph, not (-1, 1)

OpenStudy (anonymous):

ok, I get that

OpenStudy (anonymous):

So I did the diagram, so why can't 2,1 be a possible point? @ganeshie8

OpenStudy (anonymous):

thanks sir ganeshie i had almost the same problem